# Unveiling the Landscape of LLM Deployment in the Wild: An Empirical Study

### v1

@@ Abstract
Background: Large language models (LLMs) are increasingly deployed via open-source and commercial frameworks, enabling individuals and organizations to self-host advanced AI capabilities. However, insecure defaults and misconfigurations often expose LLM services to the public Internet, posing significant security and system engineering risks. Aims: This study aims to unveil the current landscape of public-facing LLM deployments in the wild through a large-scale empirical study, focusing on service prevalence, exposure characteristics, systemic vulnerabilities, and associated risks. Method: We conducted an Internet-wide measurement to identify public-facing LLM deployments across 15 frameworks, discovering 320,102 services. We extracted 158 unique API endpoints, grouped into 12 functional categories based on capabilities and security risks. We further analyzed configurations, authentication practices, and geographic distributions, revealing deployment trends and systemic issues in real-world LLM system engineering. Results: Our study shows that public LLM deployments are rapidly growing but often insecure. Among all endpoints, we observe widespread use of insecure protocols, poor TLS configurations, and unauthenticated access to critical operations. Security risks, including model disclosure, system leakage, and unauthorized access, are pervasive, highlighting the need for secure-by-default frameworks and stronger deployment practices. Conclusions: Public-facing LLM deployments suffer from widespread security and configuration flaws, exposing services to misuse, model theft, resource hijacking, and remote exploitation. Strengthening default security, deployment practices, and operational standards is critical for the growing self-hosted LLM ecosystem.
Index Terms:
large language models, LLM, LLM deployment, empirical study
@@ I Introduction
The rapid adoption of large language models (LLMs), driven by prominent models such as the GPT series from OpenAI [33] and DeepSeek’s open-source variants [1], has profoundly reshaped the landscape of artificial intelligence (AI) applications. Once confined primarily to research labs and industrial environments, these models have increasingly become accessible to the general public, fueling a widespread trend towards self-hosted and open-source deployments [18]. The availability of user-friendly tools and vibrant community ecosystems [42, 43, 38] has empowered individual enthusiasts, small enterprises, and developers to independently deploy and customize powerful language models for various personal and professional purposes, such as creative writing and content creation [21], software development and maintance [17], financial analysis and automated investment assistance [48]and personal productivity tools, greatly enriching their daily digital experiences. However, as barriers to LLM deployment continue to fall, more deployments occur without rigorous security considerations, exposing users and organizations to new operational and security risks.
@@
Motivated by these observations, we conduct a large-scale empirical analysis to systematically assess the prevalence and characteristics of public-facing LLM deployments. Our overall analysis pipeline is illustrated in Figure 1. We identify and analyze 320,102 LLM services across the Internet, spanning 15 popular deployment frameworks. From these, we extract 158 unique API endpoints, which we group into 12 functional categories based on their capabilities and associated security risks. Our study analyzes the geographic and network distribution of LLM deployments, uncovers widespread insecure configurations such as unauthenticated inference and model enumeration, and provides practical recommendations for developers and operators. The findings expose widespread security and deployment flaws, highlighting the need for secure-by-default designs and improved operational practices. Our artifacts related to this study are publicly available at https://anonymous.4open.science/r/Public-LLM-Services.
Our main contributions are as follows:
@@
  • Exposure and Security Risk Analysis: We systematically analyze 158 unique API endpoints, categorize them into 12 functional groups, and uncover systemic vulnerabilities in open-source LLM deployments, including model disclosure, system leakage, unauthorized access, vulnerabilities, and sensitive content generation.
@@
II Background and Related Work
II-A LLM Deployment Paradigms and Tooling
Deploying LLMs in real-world environments involves a multi-layered stack encompassing computation, scalable serving, user interaction, and developer support. This stack broadly consists of four key components. Inference Engines. Inference engines execute LLM computations across diverse hardware. Optimized systems like vLLM [44] and Hugging Face Transformers leverage techniques such as continuous batching, KV cache optimization, and operator fusion to improve throughput and latency [25]. Tools like LLM-Pilot [23] enable benchmarking and configuration across different environments. Model Serving Frameworks. Serving frameworks manage hosting, request routing, and scaling. Systems like Ray Serve [20] and Ollama [42] support flexible deployment across cloud and edge platforms [7, 5]. Split-based deployment further distributes model components to balance performance and privacy [6]. Application User Interfaces. User-facing applications integrate LLMs via conversational interfaces, RAG pipelines, and autonomous agents. Frameworks like LangChain and ChatGPT plugins simplify this integration but also introduce risks such as indirect prompt injection [45, 19]. Developer Tools and Ecosystems. Developer tools support fine-tuning, evaluation, and maintenance. Libraries such as DeepSpeed, Hugging Face PEFT, and LoRA enable efficient model adaptation [18, 37], while tools like TestGen-LLM [3] automate engineering workflows. Specialized accelerators are also emerging to meet large-scale inference demands [12, 7]. Despite growing modularity and accessibility, little is known about how LLM deployments are configured and exposed in practice. To bridge this gap, we conduct the first large-scale analysis of publicly accessible LLM services in the wild.
II-B Security Challenges in LLM Deployment
III Methodology
III-A Target Framework Selection
To structure our measurement and ensure broad coverage across the LLM deployment stack, we selected representative frameworks spanning four functional categories, based on the typical architecture of real-world LLM deployments: inference engines, model serving frameworks, application user interfaces (UIs), and developer tools. Inference engines (e.g., vLLM [44], llama.cpp [14]) handle local model execution optimized for performance. Model serving frameworks (e.g., Ollama [42], Ray Serve [20]) expose scalable APIs. Application UIs (e.g., Open WebUI [43], Jan [40], ComfyUI [11]) provide user-facing interfaces, while developer tools (e.g., Jupyter Notebook [22], FastAPI [36]) facilitate development workflows and integration. Frameworks were selected based on popularity in open-source communities, ease of deployment, documentation availability, and the distinctiveness of their network exposure characteristics. In total, we studied 15 widely-used frameworks, summarized in Table I.
@@ III-B Asset Discovery via FOFA
We use FOFA, a widely adopted Internet-wide asset search engine that indexes IP addresses, domains, and service metadata [9], to discover public-facing LLM deployments at scale. We constructed a feature catalog for each tool, capturing its unique network-level characteristics, to detect these deployments via FOFA. These features include default service ports (e.g., port 11434 for Ollama), page titles or HTML keywords (e.g., title=‘‘Open WebUI’’), HTTP response headers (e.g., ‘‘Ollama is running’’), and, where available, favicon hashes or known API paths, as shown in Table I. Based on this catalog, we designed custom FOFA query expressions using the platform’s advanced syntax. For instance, services like Ollama were identified using queries such as app=‘‘Ollama’’ && port=‘‘11434’’, while FastAPI-based deployments were located using expressions like title=‘‘FastAPI’’ || body=‘‘FastAPI’’) && (port=‘‘8000’’ || port=‘‘8080’’.
To improve coverage and reduce false positives, we refined these queries iteratively. This process involved testing against known deployment samples, consulting official documentation, and manually validating a representative subset of results. In particular, we examined rendered landing pages, HTTP headers, and available API metadata to confirm the identity of suspected services. As of April 20, 2025, we had identified 320,102 publicly accessible LLM-related services across 15 representative deployment tools, as shown in Table I. These include 155,423 instances of Ollama, 37,242 of Open WebUI, 28,445 of Jan, and 6,077 of vLLM, among others. These results form the foundation for the subsequent phases of our study, including endpoint probing and configuration analysis, as described in § III-C and § III-D.
@@
III-C API Endpoint Probing
We extracted 158 API endpoints from the official documentation of 15 widely used LLM deployment frameworks. These endpoints span many functionalities, including model management, file access, kernel and session management, and application deployment. As summarized in Table II, the endpoints are organized into 12 functional categories. For instance, endpoints under the Text/Chat category enable OpenAI-style text generation (e.g., /v1/chat/completions), while the Model Control group provides access to model lifecycle operations such as loading or deleting models (e.g., /models/delete). Several endpoints were found to expose potentially sensitive actions, including file uploads, session control, and even server-side code execution in certain configurations. Using these extracted paths, we constructed full URLs by combining them with the network locations of previously discovered LLM services. Probing was conducted via HTTP(S) requests, and all responses were collected and stored for further analysis. Rather than attempting a comprehensive crawl of all known services, the probing process was designed to focus on identifying insecure deployment practices across representative samples. Sampling was performed within each LLM service category to ensure statistical validity. For each category, we selected services based on a 95% confidence level and a 5% margin of error. This approach allowed us to reduce probing overhead while maintaining representative coverage across deployment types. The resulting dataset provides a comprehensive view of exposed API functionalities and their response behaviors in real-world deployments.
@@ III-D Configuration and Security Analysis
Following endpoint probing, we analyzed the returned API responses to assess the configuration and security posture of exposed LLM services. Each response was mapped to a unified schema capturing five key fields: deployment framework, endpoint category, endpoint path, response type (e.g., success, denial, error), and potential security relevance. This schema enabled consistent comparison across frameworks with heterogeneous designs and naming conventions. The analysis focused on endpoint responsiveness under unauthenticated access, functional exposure across 12 categories, and the presence of risky operations such as model management, file access, and code execution. By aggregating normalized responses across representative samples, we identified common patterns of misconfiguration, authentication gaps, and inconsistent access control enforcement.
@@ IV Results
This section presents the results of our Internet-wide empirical analysis of public LLM services. We report the overall exposure, deployment characteristics, configuration patterns, and security risks observed in the collected dataset.
IV-A General Statistics
We conduct a series of foundational statistical analyses to provide a comprehensive overview of the public landscape of self-hosted LLM services. These analyses aim to uncover the global deployment scale, organizational involvement, domain exposure, deployment surfaces, and security postures, forming a basis for deeper security investigations.
@@ IV-A1 Global and Organizational Deployment Trends
Geographic Centralization. As shown in 2(a), the United States is by far the dominant origin of public LLM services, with 111,728 instances, more than twice the count of the second-largest country, China (56,593). Other prominent contributors include Germany, Japan, and Singapore. This imbalance highlights a pronounced centralization of deployment activity within technologically advanced nations, likely driven by both cloud infrastructure maturity and early AI adoption. Notably, the “Others” category still accounts for over 53,000 instances, suggesting that self-hosted LLMs are also gaining traction in the broader global community, albeit in smaller clusters. This distribution hints at a growing democratization of LLM capabilities, though one is still heavily shaped by infrastructure access and national regulatory frameworks.
@@ IV-A2 High-Traffic Domains and Service Concentration
Beyond these unassigned cases, certain domains are associated with an unusually high number of LLM instances. As shown in Figure 3, domains such as nellasushi.es, mysuccess.be, and human-rights-law.eu each host thousands of services. This level of repetition is uncommon in conventional web deployments and suggests that these domains serve as default endpoints in automated or templated hosting environments. Two factors support this interpretation. First, services under the same domain frequently share a small number of IP addresses, indicating centralized hosting or large-scale reuse of identical deployment images. For instance, 6,206 instances under nellasushi.es are served by two IPs. Second, deployment metadata [^1] Supplemental heatmap available at https://anonymous.4open.science/r/Public-LLM-Services. reveals a consistent reliance on a limited set of frameworks such as ComfyUI, Jan, and vLLM, often in their default configurations. These patterns point to widespread use of prebuilt containers or orchestration scripts. Such concentration has both operational advantages and security implications. Focusing remediation efforts on a few high-frequency domains could reduce exposure at scale. However, any misconfiguration or compromise within these clusters could simultaneously affect thousands of endpoints. Moreover, the repeated use of domain names, IPs, and certificates erodes the reliability of trust models and complicates service attribution.
@@ IV-A5 Certificate Reuse and Identity Management
Takeaway 1: Public LLM deployments are expanding rapidly but exhibit systemic security weaknesses. Among 320,012 endpoints, over 40% use plain HTTP, more than 210,000 have missing or generic TLS metadata, and a few providers (e.g., Amazon, Cloudflare) dominate global exposure. Long-tail deployments across smaller networks add inconsistency and unpredictability. These findings highlight the urgent need for secure-by-default frameworks, better certificate management, and standardized operational practices for the self-hosted LLM ecosystem.
IV-B API Responsiveness Analysis
@@ IV-B1 Framework-Level API Responsiveness
Across exposed LLM deployments, we observe substantial variation in API responsiveness. As shown in Table IV, frameworks like Ollama and Llamafile respond to over 80% of unauthenticated API requests, suggesting permissive default configurations and minimal access control. In contrast, widely used platforms such as Open WebUI, Jan, and Text Generation WebUI exhibit responsiveness rates below 2%, indicating either frontend-only exposure or stricter backend protection. FastAPI and AnythingLLM are excluded due to the lack of standardized APIs. FastAPI serves as a generic backend framework with fully customizable endpoints, while AnythingLLM exposes inconsistent interfaces across deployments, often relying on UI interactions. Both lack the structural consistency required for comparable measurement.
Figure 6: Distribution of Reused TLS Certificate Subject CNs.
TABLE IV: Responsiveness of Different LLM Deployment Frameworks.
Framework # Resp. Sample Population Resp. Rate
Ollama 309 384 155,423 80.47%
Open WebUI 2 381 37,242 0.52%
Jan 5 380 28,445 1.32%
NextChat 1 379 25,883 0.26%
Jupyter Notebook 53 379 24,531 13.98%
ComfyUI 42 375 15,219 11.20%
Gradio 60 370 9,729 16.22%
vLLM 10 362 6,077 2.76%
llama.cpp 36 353 4,234 10.20%
GPT4All 13 335 2,572 3.88%
Text Generation Web UI 2 324 2,051 0.62%
Ray Serve 1 188 365 0.53%
Llamafile 29 36 39 80.56%
@@
The HTTP status codes returned by these services reflect their underlying security posture. While some services return 200, exposing full functionality, many respond with 401 (unauthorized), 403 (forbidden), or 404 (not found), indicating authentication checks, access restrictions, or endpoint obfuscation. A smaller number return 5xx errors, suggesting unstable or misconfigured services. We also observe a notable number of 400 Bad Request responses, mainly from Gradio-based deployments. These indicate that the service is reachable but the API is undocumented, unsupported, or not intended for direct programmatic use, consistent with Gradio’s focus on interactive UIs rather than formal APIs. These response patterns reveal implicit security mechanisms across frameworks. High responsiveness is often associated with frameworks that expose APIs by default and lack built-in protection, whereas lower responsiveness generally corresponds to more defensive configurations. Notably, a non-trivial portion of services respond with structured denial (401/403), suggesting that some frameworks implement basic safeguards even when deployed in public environments.

### v2

@@ Abstract
Large language models (LLMs) are increasingly deployed through open-source and commercial frameworks, enabling individuals and organizations to self-host advanced LLM capabilities. As LLM deployments become prevalent, particularly in industry, ensuring their secure and reliable operation has become a critical issue. However, insecure defaults and misconfigurations often expose LLM services to the public internet, posing serious security and system engineering risks. This study conducted a large-scale empirical investigation of public-facing LLM deployments, focusing on the prevalence of services, exposure characteristics, systemic vulnerabilities, and associated risks. Through internet-wide measurements, we identified 320,102 public-facing LLM services across 15 frameworks and extracted 158 unique API endpoints, categorized into 12 functional groups based on functionality and security risk. Our analysis found that over 40% of endpoints used plain HTTP, and over 210,000 endpoints lacked valid TLS metadata. API exposure was highly inconsistent: some frameworks, such as Ollama, responded to over 35% of unauthenticated API requests, with about 15% leaking model or system information, while other frameworks implemented stricter controls. We observed widespread use of insecure protocols, poor TLS configurations, and unauthenticated access to critical operations. These security risks, such as model leakage, system compromise, and unauthorized access, are pervasive and highlight the need for a secure-by-default framework and stronger deployment practices.
@@ I Introduction
Driven by renowned models like OpenAI’s GPT series [33] and DeepSeek’s open-source variant [9], large language models (LLMs) are rapidly gaining popularity and profoundly reshaping a wide range of applications. Once primarily confined to research labs and industrial environments, these models are now not only continuously deployed in-depth within the industry, but are also gradually opened to the wider public, promoting the vigorous development of self-hosted and open source deployment. The emergence of user-friendly tools and a vibrant community ecosystem [42, 43, 38] has enabled individual enthusiasts, small enterprises, and developers to independently deploy and customize powerful LLMs for a variety of personal and professional needs, such as creative writing and content creation [20], software development and maintenance [16], financial analysis and automated investment assistance [48], and personal productivity tools, significantly enriching their daily digital experiences. However, as the barrier to entry for LLM deployment continues to decrease, an increasing number of deployments lack rigorous security considerations, exposing users and organizations to new operational and security risks.
@@
Motivated by these observations, we conduct a large-scale empirical study to examine the prevalence and characteristics of public-facing LLM deployments. Our investigation is guided by the following research questions (RQs):
RQ1 What is the global distribution and typical setup of public LLM services? We aim to map the current landscape of public LLM deployments, including their geographic, organizational, and technical characteristics.
RQ2 How do major LLM deployment frameworks differ in API exposure and access control? This helps us understand the diversity of deployment strategies and the effectiveness of default security mechanisms.
RQ3 What are the main security risks and vulnerabilities in public LLM deployments? We seek to identify prevalent threats and systemic issues that may affect the security of public LLM services.
To answer these RQs, our analysis pipeline is shown in Figure 1. We identify and analyze 320,102 public LLM services, spanning 15 popular deployment frameworks. From these, we extract 158 unique API endpoints, which we group into 12 functional categories based on their capabilities and associated security risks. Our study analyzes the geographic and network distribution of LLM deployments, uncovers widespread insecure configurations such as unauthenticated inference and model enumeration, and provides practical recommendations for developers and operators. The findings expose widespread security and deployment flaws, highlighting the need for secure-by-default designs and improved operational practices.
Contributions. Our primary contributions [^1] Our artifacts are publicly available at https://github.com/cccccindy19/LLM-Services. are as follows:
@@
  • Exposure and Security Risk Analysis: We systematically analyze 158 unique API endpoints, categorize them into 12 functional groups, and uncover systemic vulnerabilities in open-source LLM deployments.
@@
II Background and Related Work
II-A LLM Deployment Paradigms and Tooling
An LLM framework provides the necessary foundation for building, deploying, and operating LLM-powered services in real-world environments. At a high level, an LLM deployment framework is a software platform or toolkit that enables organizations to efficiently and securely operate LLMs and connect models with users and applications at scale. Such frameworks typically integrate multiple layers of computing, service management, user interaction, and developer support.
This stack broadly consists of four key components. Inference Engines. Inference engines execute LLM computations across diverse hardware. Optimized systems like vLLM [44] and Hugging Face Transformers leverage techniques such as continuous batching, KV cache optimization, and operator fusion to improve throughput and latency [24]. Tools like LLM-Pilot [22] support benchmarking and tuning in production environments. Model Serving Frameworks. Serving frameworks manage hosting, request routing, and scaling. Systems like Ray Serve [19] and Ollama [42] support flexible deployment across cloud and edge platforms [5, 3]. Split-based deployment further distributes model components to balance performance and privacy [4]. Application User Interfaces. User-facing applications integrate LLMs via conversational interfaces, RAG pipelines, and autonomous agents. Frameworks like LangChain and ChatGPT plugins simplify this integration but also introduce risks such as indirect prompt injection [45, 18]. Developer Tools and Ecosystems. Developer tools support fine-tuning, evaluation, and maintenance. Libraries such as DeepSpeed, Hugging Face PEFT, and LoRA enable efficient model adaptation [17, 37], while tools like TestGen-LLM [1] automate engineering workflows. Specialized accelerators are also emerging to meet large-scale inference demands [11, 5]. Despite growing modularity and accessibility, little is known about how LLM deployments are configured and exposed in practice. To bridge this gap, we conduct the first large-scale analysis of publicly accessible LLM services in the wild.
II-B Security Challenges in LLM Services
III Methodology
III-A Target Framework Selection
To structure our measurement and ensure broad coverage across the LLM deployment stack, we selected representative frameworks across four functional categories: Inference engines (e.g., vLLM [44], llama.cpp [13]) handle local model execution optimized for performance. Model serving frameworks (e.g., Ollama [42], Ray Serve [19]) expose scalable APIs. Application UIs (e.g., Open WebUI [43], Jan [40], ComfyUI [10]) provide user-facing interfaces, while developer tools (e.g., Jupyter Notebook [21], FastAPI [36]) facilitate development workflows and integration. Our framework selection followed a multi-step process: First, we surveyed the most cited projects in open source communities (e.g., GitHub, Hugging Face) and industry forums to compile an initial shortlist. We then screened these frameworks based on factors such as ease of deployment (e.g., availability of Docker images or installation scripts), quality and completeness of documentation, and evidence of active maintenance (recent commits, issue activity). Next, we prioritized frameworks that exhibited unique network exposure patterns and supported diverse deployment scenarios, such as on-premises, cloud, and hybrid deployments. This process enabled us to capture both widely adopted solutions and emerging tools with unique characteristics. In total, we selected 15 widely used frameworks, which are summarized in Table I.
@@ III-B Asset Discovery via FOFA
We use FOFA, a widely adopted Internet-wide asset search engine that indexes IP addresses, domains, and service metadata [7], to discover public-facing LLM deployments at scale. For each tool, we built a feature catalog that captures its unique network-level characteristics, such as default port (e.g., Ollama’s default port is 11434), page title or HTML keywords (e.g., title=Open WebUI’’), HTTP headers (e.g., Ollama is running’’), favicon hashes, and known API paths (see Table I for details). Based on this catalog, we designed custom FOFA query expressions using the platform’s advanced syntax. For instance, app=Ollama’’ && port=11434’’ for Ollama and title=FastAPI’’ || body=FastAPI’’ && (port=8000’’ || port=8080’’) for FastAPI). To improve coverage and reduce false positives, we refined these queries iteratively. This process involved testing against known deployment samples, consulting official documentation, and manually validating a representative subset of results. In particular, we examined rendered landing pages, HTTP headers, and available API metadata to confirm the identity of suspected services. As of April 20, 2025, we had identified 320,102 publicly accessible LLM-related services across 15 representative deployment tools, as shown in Table I. These results form the foundation for the subsequent phases of our study, including endpoint probing and configuration analysis.
III-C API Endpoint Probing
We extracted 158 API endpoints from the official documentation of 15 widely used LLM deployment frameworks. These endpoints span many functionalities, including model management, file access, kernel and session management, and application deployment. As summarized in Table II, the endpoints are organized into 12 functional categories. For instance, endpoints under the Text/Chat category enable OpenAI-style text generation (e.g., /v1/chat/completions), while the Model Control group provides access to model lifecycle operations such as loading or deleting models (e.g., /models/delete). Several endpoints were found to expose potentially sensitive actions, including file uploads, session control, and even server-side code execution in certain configurations. Using the extracted paths, we combined them with the network locations of previously identified LLM services to construct complete URLs. An automated process then probed these URLs with HTTP(S) requests, systematically collecting and storing all responses for further analysis. The resulting dataset provides a comprehensive view of exposed API functionalities and their response behaviors in real-world deployments.
@@ III-D Configuration and Security Analysis
Following endpoint probing, we analyzed the returned API responses to assess the configuration and security posture of exposed LLM services. Each response was mapped to a unified schema capturing five key fields: deployment framework, endpoint category, endpoint path, response type (e.g., success, denial, error), and potential security relevance. This schema enabled consistent comparison across frameworks with heterogeneous designs and naming conventions. The analysis focused on endpoint responsiveness under unauthenticated access, functional exposure across 12 categories, and the presence of risky operations such as model management, file access, and code execution.
@@ IV Results
This section presents the results of our empirical analysis of public LLM services. Findings are organized by RQs, covering overall exposure, deployment characteristics, configuration patterns, and security risks observed in the dataset.
IV-A RQ1: General Statistics
We begin with basic statistical analyses to summarize the global deployment, organizational involvement, exposure surfaces, and security postures of self-hosted LLM services.
@@ IV-A1 Global and Organizational Deployment Trends
Geographic Centralization. As shown in 2(a) , the United States leads with 111,728 public LLM services, more than double the number of second-place China (56,593). Other prominent contributors include Germany, Japan, and Singapore. This imbalance highlights a pronounced centralization of deployment activity within technologically advanced nations, likely driven by both cloud infrastructure maturity and early AI adoption. Notably, the “Others” category still accounts for over 53,000 instances, suggesting that self-hosted LLMs are also gaining traction in the broader global community, albeit in smaller clusters. This distribution hints at a growing democratization of LLM capabilities, though one is still heavily shaped by infrastructure access and national regulatory frameworks.
@@ IV-A2 High-Traffic Domains and Service Concentration
Beyond these unassigned cases, certain domains are associated with an unusually high number of LLM instances. As shown in Figure 3, domains such as nellasushi.es, mysuccess.be, and human-rights-law.eu each host thousands of services. This level of repetition is uncommon in conventional web deployments and suggests that these domains serve as default endpoints in automated or templated hosting environments. Two factors support this interpretation. First, services under the same domain frequently share a small number of IP addresses, indicating centralized hosting or large-scale reuse of identical deployment images. For instance, 6,206 instances under nellasushi.es are served by two IPs. Second, deployment metadata [^2] Supplemental heatmap available at https://github.com/cccccindy19/LLM-Services. reveals a consistent reliance on a limited set of frameworks such as ComfyUI, Jan, and vLLM, often in their default configurations. These patterns point to widespread use of prebuilt containers or orchestration scripts. Such concentration has both operational advantages and security implications. Focusing remediation efforts on a few high-frequency domains could reduce exposure at scale. However, any misconfiguration or compromise within these clusters could simultaneously affect thousands of endpoints. Moreover, the repeated use of domain names, IPs, and certificates erodes the reliability of trust models and complicates service attribution.
@@ IV-A5 Certificate Reuse and Identity Management
Answer to RQ1: Public LLM deployments are spread across the globe, with most exposure concentrated on a few major providers (e.g., Amazon, Cloudflare), but also encompassing many smaller networks. Of the 320,012 endpoints, over 40% use plain HTTP, and over 210,000 lack valid TLS metadata. Deployment setup and security practices are highly inconsistent, highlighting the need for standardization and secure-by-default practices.
Figure 6: Distribution of Reused TLS Certificate Subject CNs.
IV-B RQ2: API Exposure
@@ IV-B1 Framework-Level API Responsiveness
We observed significant variation in API responsiveness among publicly available LLM deployments. As shown in Table IV, frameworks such as Ollama and Llamafile responded to over 10% of unauthenticated API requests, indicating permissive defaults and minimal access controls. In contrast, platforms such as Open WebUI, NextChat, and Gradio had response rates below 0.1%, suggesting more stringent backend protection or exposing only the frontend. FastAPI and AnythingLLM were excluded due to their lack of standardized and comparable APIs. Due to the large number of Ollama deployments, we conducted API probing on a small random sample of 877 instances. As shown in the first row of Table IV , 35.23% of Ollama instances responded. However, upon closer inspection, we discovered that a significant portion of these responses consisted of cookie-cutter “template responses” (generic, uninformative replies) rather than meaningful API output. We speculate that these template responses may be due to default configurations, placeholder endpoints, or defensive mechanisms designed to prevent automated probing. Excluding template responses reduced the actual success rate to 15.28%, closely matching the overall response rate (14.69%) observed in the full Ollama dataset of 155,423 instances. Notably, we did not observe such template response patterns in other frameworks.
TABLE IV: Responsiveness of Deployment Frameworks.
Framework # Resp. Sample / Total Resp. Rate
Ollama 309 (w/)1 877 35.23%
134 (w/o)2 877 15.28%
Ollama 22,831 155,423 14.69%
Open WebUI 10 43,947 0.02%
Jan 7 380 1.84%
NextChat 13 20,000 0.07%
Jupyter Notebook 1,626 24,532 6.63%
ComfyUI 357 6,241 5.72%
Gradio 7 12,727 0.06%
vLLM 3 948 0.32%
llama.cpp 110 1,460 7.53%
GPT4All 23 732 3.14%
Text Generation Web UI 1 324 0.31%
Ray Serve 2 188 1.06%
Llamafile 42 412 10.19%
  • 1
    With (w/) template response.
  • 2
    Without (w/o) template response.
@@
The returned HTTP status codes also reflected security: some services returned 200 (full access), while many returned 401, 403, or 404, indicating authentication, access control, or endpoint obfuscation. Gradio-based deployments often returned 400, reflecting that the API was not designed for direct programmatic use. These response patterns reveal implicit security mechanisms across frameworks. High response rates are often associated with frameworks that expose APIs by default and lack built-in protection, while lower response rates generally correspond to more defensive configurations. It is worth noting that a significant portion of services respond with structured rejections (401/403), which suggests that some frameworks implement basic security measures even when deployed in public environments.